# "Which LLM should I use?": Evaluating LLMs for tasks performed by Undergraduate Computer Science Students in India

### v1

@@ 5. Discussion
As shown in the results, no LLM performs best in every task, and has strengths and weaknesses. We suggest students and teachers can choose the best LLM based on their varying needs and not use any LLM blindly. For content writing purposes, it is advisable to initially create a rudimentary summary of the requirements and subsequently request the LLM to expand upon it. This methodology can aid in minimizing the likelihood of errors in the response. LLMs can function as a proofreading tool in writing tasks. For technical tasks, Nevertheless, students should be cognizant that LLMs may encounter difficulties with tasks necessitating mathematical and logical reasoning and may provide incorrect answers. Instead of attempting to generate complete code, students can employ LLMs to comprehend different algorithms and codebases, learn new frameworks and concepts, and suggest possible enhancements to the code. LLMs can also generate sample test cases, and help to clarify any doubt about a potential solution. However, students should be cautious about potential plagiarism and are advised to always acknowledge their sources and decide the usage of LLMs depending on the specific policy implemented by the instructor.
Instructors can utilize LLMs to design questions that promote critical thinking and logical reasoning skills, which are not directly solvable by LLMs as shown by our results. This can help enhance students’ cognitive abilities. Additionally, LLMs can be used to provide personalized questions to cater to the varying learning needs of students. In order to discourage plagiarism, instructors can give specific and detailed questions that require students to cite class content and provide personal reflections on resources not commonly found on the web, such as critiques on research papers. At the same time, students can be encouraged to use LLMs to learn how to write critiques, read research papers, and research related topics. LLMs can then be used to provide feedback to students, enabling them to improve their work further. In some courses, instructors can also design assignments that require students to develop real-world applications with specific requirements, which necessitate collaboration and an in-depth understanding of multiple concepts. Students can be asked to justify their design choices and can use LLMs for ideation, brainstorming, and learning. Instructors and universities can collaborate with researchers working on large language models to fine-tune LLMs specifically for the education domain. Microsoft’s AdaptLLM (Cheng et al., 2023) has shown promising results in this area.
@@ 5.1. Limitations and Future Work
It is evident that LLMs are continuously being updated with better models. It remains to be seen when these models will overcome the limitations presented in this paper. With new LLMs such as Google Gemini, Mistral’s Mixtral(8x7B) and Medium coming up, we plan to include these LLMs in our evaluation in the future. We were ourselves responsible for creating most of the datasets used in this study’s evaluation which may have also added some undesirable bias in this study. We also acknowledge that the results and recommendations made in this paper are based on our own interpretations. To further establish their validity, we plan to evaluate these recommendations using a larger study.
@@ 6. CONCLUSION
In this paper, we analyzed the strengths and weaknesses of ChatGPT, Microsoft Copilot, GitHub Copilot Chat, and Google Bard on diverse tasks relevant in undergraduate computer science education. We find that different LLMs are best-suited for different tasks and there is no one LLM performing well on all the tasks. This study can be helpful for both students and instructors in selecting a specific LLM for any task. We concluded with some recommendations and suggestions for both students and instructors.

### v2

@@ 5. DISCUSSION
In our study, we conducted a comprehensive evaluation of various large language models (LLMs) frequently utilized by undergraduate computer science students, including Google Bard, ChatGPT(3.5), GitHub Copilot Chat, and Microsoft Copilot. Our findings reveal distinctive strengths across different tasks. ChatGPT and GitHub Copilot Chat exhibited remarkable proficiency in solving Leetcode.com questions across all difficulty levels. Microsoft Copilot showcased exceptional performance in providing code explanations and documentation and excelled in theoretical and humanities assignments as well. Additionally, GitHub Copilot Chat emerged as the top performer for programming assignments, while Google Bard proved to be the optimal choice for learning new concepts and frameworks. ChatGPT demonstrated the best performance in composing emails.
In this study, we anticipate that LLMs will be of significant use to students and instructors in the near future. Hence for them, it becomes increasingly important to know which LLM performs best for a given task. Our findings show that no LLM works best across all the tasks. This paper thus provides a much-needed starting point underlining the strengths and weaknesses of each LLM. We encourage students and instructors to use this study’s insights along with their own assessments to make a well-informed choice best suited to their tasks and needs.
Below, we present the recommendations for students and instructors to effectively integrate LLMs into their workflow. Wherever possible, we also cite existing work where some of these recommendations have also been proposed.
5.1. Recommendation for students:
  • Content Writing: When utilizing LLMs for writing tasks such as email composition, writing cover letters, etc, it is advisable to initially create a rudimentary summary of the requirements and subsequently request the LLM to expand upon it. This methodology can help to address the potential accuracy and reliability concerns raised by (Budhiraja et al., 2024; Joshi et al., 2023b). They can also assist in proofreading content. Nevertheless, students are advised to consistently verify the response for any potential inaccuracies, and acknowledge their sources.
  • Academic Tasks: LLMs can serve as an ideation and research instrument for academic tasks such as assignment completion. Instead of directly solving the assignment, they can assist students in selecting a project idea(Joshi et al., 2023a), providing a fundamental outline and plan of action for the project, and comprehending problems from diverse perspectives. This could help mitigate the concerns highlighted by (Lau and Guo, 2023), which may adversely affect student learning outcomes.
  • Computer Science Tasks: LLMs can be advantageous in selecting between various frameworks and technologies by presenting the pros and cons of each option(Daun and Brings, 2023). Nevertheless, students should be cognizant that LLMs may encounter difficulties with tasks necessitating logical reasoning and may provide incorrect answers as shown in our findings and (Sarsa et al., 2022; Balse et al., 2023; Budhiraja et al., 2024). Instead of attempting to generate complete code, students can employ LLMs to generate code explanations(Leinonen et al., 2023; MacNeil et al., 2023; Sarsa et al., 2022; Wermelinger, 2023), learn new frameworks and concepts, and suggest possible enhancements to the code. LLMs can also generate sample test cases(Joshi et al., 2023a), and help to clarify any doubt about a potential solution.
5.2. Recommendation for Instructors
  • Instructors can provide guidance for using LLMs using insights shared by this study and various other guiding studies(Joshi et al., 2023b, a; Daun and Brings, 2023; Lau and Guo, 2023; Budhiraja et al., 2024; MacNeil et al., 2023). They can try to integrate them into their courses rather than accepting the unsupervised use by students, which can negatively impact their learning(Daun and Brings, 2023).
  • Instructors can utilize LLMs to design questions that promote critical thinking and logical reasoning skills, which are not directly solvable by LLMs as suggested by (Budhiraja et al., 2024; Joshi et al., 2023a) and our results. This can help enhance students’ cognitive abilities. Additionally, LLMs can be used to provide personalized questions to cater to the varying learning needs of students(Daun and Brings, 2023).
  • Social Sciences and Humanities (SSH) assignments, are more prone to plagiarism by students as LLMs particularly excel in content and essay writing(Herbold et al., 2023). Instructors can give specific and detailed questions that require students to cite class content and provide personal reflections on resources not commonly found on the web, such as critiques of research papers. Here, students can be encouraged to use LLMs to learn how to write critiques, read research papers, and research related topics. LLMs can then be used to provide feedback to students, enabling them to improve their work further(Joshi et al., 2023a; Daun and Brings, 2023).
  • In computer science subjects, instructors can design assignments that require students to develop real-world applications with specific requirements(Joshi et al., 2023a; Daun and Brings, 2023; Lau and Guo, 2023), which necessitate collaboration and an in-depth understanding of multiple concepts. Students can be asked to justify their design choices and can use LLMs for ideation, brainstorming, and learning.
  • Instructors can explore new modes of open-ended evaluations, such as giving students a topic and asking them to design a problem around it, including solutions and test cases (if applicable) (Joshi et al., 2023a).
  • Instructors and universities can collaborate with researchers working on large language models to develop LLMs specifically for the education domain. Work done by Microsoft(Yee and Yee, 2024) has demonstrated remarkable progress in this step.
@@ 5.3. Limitations and Future Work
Significant advancements are happening in research related to Large Language Models and new models such as Anthropic’s Claude 3, Mistral’s Mixtral(8x7B), Medium and Large have been announced. The existing tools such as Google Bard(Pichai, 2023b), and Microsoft Copilot(Gedeon, 2024) are being updated with better models. However, Chen et al(Chen et al., 2023) suggests that LLMs quality is also taking a hit with time. We plan to incorporate these advancements in our evaluation in the future.
One of the limitations of this study is the diversity of tasks covered, as we focused on a specific set of tasks commonly encountered by undergraduate computer science students in India. This may not fully capture the potential effectiveness of LLMs across a broader range of tasks relevant to undergraduate computer science education.
We were ourselves responsible for creating most of the datasets used in this study’s evaluation which may have also added some undesirable bias to this study. We also acknowledge that the results and recommendations made in this paper are based on our own interpretations. We plan to evaluate these recommendations using a larger study to further establish their validity.
@@ 6. CONCLUSION
This study evaluated the strengths and weaknesses of ChatGPT(3.5), Microsoft Copilot, GitHub Copilot Chat, and Google Bard across various tasks relevant to undergraduate computer science education. The results indicated that different large language models (LLMs) excel at different tasks, with no single LLM outperforming the others across all tasks. These findings can guide students and instructors in selecting the most appropriate LLM for a given educational task. The study concludes by offering recommendations for students and instructors on effectively leveraging LLMs in computer science education to enhance learning outcomes.